# ALINE: Joint Amortization for Bayesian Inference and Active Data Acquisition

### v1

@@ 3.4 Architecture
Aline employs a single, integrated neural architecture based on Transformer Neural Process (TNP) [50, 13], to concurrently manage historical observations, propose future queries, and condition predictions on specific, potentially varying, inference objectives. An overview of Aline’s architecture is provided in Figure A1.
The model inputs are structured into three sets. Following standard TNP-based architecture, the context set $\mathcal{D}_{t}=\{(x_{i},y_{i})\}_{i=1}^{t}$ comprises the history of observations, and the target set $\mathcal{T}$ contains the specific target specifier $\xi$. To facilitate active data acquisition, we incorporate a query set $\mathcal{Q}=\{x^{\text{q}}_{n}\}_{n=1}^{N}$ of candidate points. In this paper, we focus on a discrete pool-based setting for consistency, though it can be straightforwardly extended to continuous design spaces (e.g., [61]). Details regarding the embeddings of these inputs are provided in Section B.2.
Standard transformer attention mechanisms process these embedded representations. Self-attention operates within the context set, capturing dependencies within $\mathcal{D}_{t}$. Both the query and target set then employ cross-attention to attend to the processed context set representations. To enable the policy $\pi_{\psi}$ to dynamically adapt its acquisition strategy based on the specific inference goal $\xi$, we introduce an additional query-target cross-attention mechanism to allow the query candidates to directly attend to the target set. This allows the evaluation of each candidate $x^{\text{q}}_{n}$ to be informed by its relevance to the different potential targets $\xi$. Examples of the attention mask are shown in Figure A2.
Finally, two specialized output heads operate on these processed representations. The inference head ($q_{\phi}$), following ACE [13], uses a Gaussian mixture to parameterize the approximate posteriors and posterior predictives. The acquisition head ($\pi_{\psi}$) generates a policy over the query set $\pi_{\psi}(x_{t+1}|\mathcal{D}_{t})$, drawing on principles from policy-based design methods [34, 47]. This novel combination within a TNP framework allows Aline to effectively integrate established components for inference and acquisition with a new mechanism for adaptive goal-oriented querying.
4 Related work
A major family of ABI methods is Neural Processes (NPs) [25, 24], which learn a mapping from observed context data points to a predictive distribution for new target points. Early NPs often employed MLP-based encoders [25, 24, 35, 9], while more recent works utilize more advanced attention and transformer architectures [39, 50, 49, 17, 18, 3, 2]. Complementary to NPs, methods within simulation-based inference [15] focus on amortizing the posterior distribution [51, 46, 32, 57, 48, 71]. More recently, methods for amortizing both the posterior and posterior predictive distributions have been proposed [29, 56, 13]. Specifically, ACE [13] shows how to flexibly condition on diverse user-specified inference targets, a method Aline incorporates for its own flexible inference capabilities. Building on this principle of goal-directed adaptability, Aline advances it by integrating a learned policy that dynamically tailors the data acquisition strategy to such specified objectives.
@@ 4 Related work
Existing amortized BED or BAL methods [21, 37, 7, 43] that learn an offline design policy do not provide real-time estimates of the posterior, unlike Aline. An exception to this is the recent work of Blau et al., [6], which uses a variational posterior bound to provide amortized posterior inference via a separate proposal network. Compared to their method, Aline uses a single, unified architecture where the same model performs amortized inference for both posterior and posterior predictive distributions, and learns the flexible acquisition policy. Recently, transformer-based models have been explored in various amortized sequential decision-making settings, including decision-aware BED [34], Bayesian optimization (BO) [47], multi-objective BO [36], and preferential BO [72].
@@ 5 Experiments
We now empirically evaluate Aline’s performance in different active data acquisition and amortized inference tasks. We begin with the active learning task in Section 5.1, where we want to efficiently minimize the uncertainty over an unknown function by querying $T$ data points. Then, we test Aline’s policy on standard BED benchmarks in Section 5.2. Finally, in Section 5.3, we demonstrate the benefit of Aline’s flexible targeting feature in a psychometric modeling task [68]. The code to reproduce our experiments is available at: https://github.com/huangdaolang/aline.
@@ 5.1 Active learning for regression and hyperparameter inference
For the active learning task, Aline is trained on a diverse collection of fully synthetic functions drawn from Gaussian Process (GP) [59] priors (see Section C.1.1 for details). We evaluate Aline’s performance under both in-distribution and out-of-distribution settings. For the in-distribution setting, Aline is evaluated on synthetic functions sampled from the same GP prior that is used during training. In the out-of-distribution setting, we evaluate Aline on benchmark functions (Higdon, Goldstein-Price, Ackley, and Gramacy) unseen during training, to assess generalization beyond the training regime. We compare Aline against non-amortized GP models equipped with standard acquisition functions such as Uncertainty Sampling (GP-US), Variance Reduction (GP-VR) [70], EPIG (GP-EPIG) [63], and Random Sampling (GP-RS). Additionally, we include an amortized inference baseline, the Amortized Conditioning Engine (ACE) [13], paired with Uncertainty Sampling (ACE-US), to specifically evaluate the advantage of Aline’s learned acquisition policy over using a standard acquisition function with an amortized inference model. The performance metric is the root mean squared error (RMSE) of the predictions on a held-out test set.
@@
Results for the active learning task in Figure 2 show that Aline performs comparably to the best-performing GP-based methods for the in-distribution setting. Importantly, for the out-of-distribution setting, Aline outperforms the baselines in 3 out of the 4 benchmark functions. These results highlight the advantage of Aline’s end-to-end learning strategy, which obviates the need for kernel specification using GPs or explicit acquisition function selection. Further evaluations on additional benchmark functions (Gramacy 1D, Branin, Three Hump Camel), visualizations of Aline’s sequential querying strategy with corresponding predictive updates, and a comparison of average inference times are provided in Section D.1.
Additionally, for the in-distribution setting, we test Aline’s capability to infer the underlying GP’s hyperparameters—without retraining, by leveraging Aline’s flexible target specification at runtime. For baselines, we use ACE-US and ACE-RS, since ACE [13] is also capable of posterior estimation. Figure 3 shows that Aline yields higher log probabilities of the true parameter value under the estimated posterior at each step compared to the baselines. This is due to the ability to flexibly switch Aline’s acquisition strategy to parameter inference, unlike other active learning methods. We also visualize the obtained posteriors in Section D.1.
@@ 5.2 Benchmarking on Bayesian experimental design tasks
We test Aline on two classical BED tasks: Location Finding [62] and Constant Elasticity of Substitution (CES) [1], with two- and six-dimensional design space, respectively. As baselines, we include a random design policy, a gradient-based method with variational Prior Contrastive Estimation (VPCE) [23], and two amortized BED methods: Deep Adaptive Design (DAD) [21] and RL-BOED [7]. Details of the tasks and the baselines are provided in Section C.2.
To evaluate performance, we compute a lower bound of the total EIG, namely the sequential Prior Contrastive Estimation lower bound [21]. As shown in Table 2, Aline surpasses all the baselines in the Location Finding task and achieves performance comparable to RL-BOED in the CES task. Notably, Aline’s training time is reduced as its reward is based on the internal posterior improvement and does not require a large number of contrastive samples to estimate sEIG. While Aline’s deployment time is slightly higher than MLP-based amortized methods due to the computational cost of its transformer architecture, it remains orders of magnitude faster than non-amortized approaches like VPCE. Visualizations of Aline’s inferred posterior distributions are provided in Section D.2.
@@
EIG lower bound ($\uparrow$) Training time (h) Deployment time (s) EIG lower bound ($\uparrow$) Training time (h) Deployment time (s)
Random 5.17$\pm$0.05 N/A N/A 9.05$\pm$0.26 N/A N/A
VPCE [23] 5.25$\pm$0.22 N/A 146.59$\pm$0.09 9.40$\pm$0.27 N/A 788.90$\pm$1.03
DAD [21] 7.33$\pm$0.06 7.24 0.0001$\pm$0.00 10.77$\pm$0.15 13.70 0.0001$\pm$0.00
RL-BOED [7] 7.70$\pm$0.06 63.29 0.0003$\pm$0.00 14.60$\pm$0.10 67.28 0.0004$\pm$0.00
Aline (ours) 8.91$\pm$0.04 21.20 0.03$\pm$0.00 14.37$\pm$0.08 13.29 0.04$\pm$0.00
@@ 5.3 Psychometric model
Our final experiment involves the psychometric modeling task [68]—a fundamental scenario in behavioral sciences, from neuroscience to clinical settings [27, 54, 69], where the goal is to infer parameters governing an observer’s responses to varying stimulus intensities. The psychometric function used here is characterized by four parameters: threshold, slope, guess rate, and lapse rate; see Section C.3.1 for details. Different research questions in psychophysics necessitate focusing on different parameter subsets. For instance, studies on perceptual sensitivity primarily target precise estimation of threshold and slope, while investigations into response biases or attentional phenomena might focus on the guess and lapse rates. This is where Aline’s unique flexible querying strategy can be used to target specific parameter subsets of interest.
We compare Aline with two established adaptive psychophysical methods: QUEST+ [66], which targets all parameters simultaneously, and Psi-marginal [55], which can marginalize over nuisance parameters to focus on a specified subset, a non-amortized gold-standard method for flexible acquisition. We evaluate scenarios targeting either the threshold and slope parameters or the guess and lapse rates. Details of the baselines and the experimental setup are in Section C.3.2.
@@
Figure 4 shows the results. When targeting threshold and slope (Figure 4a), which are generally easier to infer, Aline achieves results comparable to baselines. When targeting guess and lapse rates (Figure 4c), QUEST+ performs sub-optimally as its experimental design strategy is dominated by the more readily estimable threshold and slope parameters. In contrast, both Psi-marginal and Aline lead to significantly better inference than QUEST+ when explicitly targeting guess and lapse rates. Moreover, Aline offers a 10×10$\times$10 × speedup over these non-amortized methods (See Appendix D). We also visualize the query strategies adopted by Aline in the two cases: when targeting threshold and slope (Figure 4b), stimuli are concentrated near the estimated threshold. Conversely, when targeting guess and lapse rates (Figure 4d), Aline appropriately selects ‘easy’ stimuli at extreme values where mistakes can be more readily attributed to random behavior (governed by lapses and guesses) rather than the discriminative ability of the subject (governed by threshold and slope).
@@ 6 Conclusion
Currently, Aline operates with pre-defined, fixed priors, necessitating re-training for different prior specifications. Future work could explore prior amortization [13, 67], to allow for dynamic prior conditioning. As Aline estimates marginal posteriors, extending this to joint posterior estimation, potentially via autoregressive modeling [9], is a promising direction. Note that, at deployment, we may encounter observations that differ substantially from the training data, leading to degradation in performance. This issue can potentially be tackled by combining Aline with robust approaches such as [20, 33]. Lastly, Aline’s current architecture is tailored to fixed input dimensionalities and discrete design spaces, a common practice with TNPs [50, 47, 72, 34]. Generalizing Aline to be dimension-agnostic [42] and to support continuous experimental designs [61] are valuable avenues for future research.
@@ Acknowledgements
DH, LA and SK were supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI). LA was also supported by Research Council of Finland grants 358980 and 356498. SK was also supported by the UKRI Turing AI World-Leading Researcher Fellowship, [EP/W002973/1]. AB was supported by the Research Council of Finland grant no. 362534. The authors wish to thank Aalto Science-IT project, and CSC–IT Center for Science, Finland, for the computational and data storage resources provided.
@@
  • In Appendix A, we provide detailed derivations and proofs for the theoretical claims made regarding information gain and variational bounds.
  • In Appendix B, we present the complete training algorithm and the specifics of the Aline model.
  • In Appendix C, we provide comprehensive details for each experimental setup, including task descriptions and baseline implementations.
  • In Appendix D, we present additional experimental results, including further visualizations, performance on more benchmarks, and analyses of inference times.
  • In Appendix E, we provide an overview of the computational resources and software dependencies for this work.
@@ A.2 Proof of Proposition 1
$\text{sEPIG}(\psi)$ $:=\mathbb{E}_{p_{\star}(x^{\star})p(\mathcal{D}\,|\,\pi_{\psi})}[\text{H}[p({y}^{\star}\,|\,x^{\star})]-\text{H}[p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})]]$
@@
[…]
@@ A.3 Proof of Proposition 2
Using the expressions for $\text{sEIG}_{\theta_{S}}$ and $\mathcal{J}_{S}^{\theta}$ from Eq. 7 and Eq. 8, respectively, and noting that $q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})=\sum_{l\in S}q_{\phi}(\theta_{l}\,|\,\mathcal{D}_{T})$, we can write the expression for $\text{sEIG}_{\theta_{S}}(\psi)-\mathcal{J}_{S}^{\theta}(\psi)$ as:
@@
$=\mathbb{E}_{p(\mathcal{D}_{T},\theta\,|\,\pi_{\psi},\theta)}\left[\log\frac{p(\theta_{S}\,|\,\mathcal{D}_{T})}{q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})}\right]$ (A8)
$=\mathbb{E}_{p(\mathcal{D}_{T},\theta_{S}\,|\,\pi_{\psi},\theta)}\left[\log\frac{p(\theta_{S}\,|\,\mathcal{D}_{T})}{q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})}\right]$ (A9)
@@
Here, Eq. A8 follows from the fact $p(\theta)p(\mathcal{D}_{T}\,|\,\pi_{\psi},\theta)=p(\mathcal{D}_{T},\theta\,|\,\pi_{\psi},\theta)$, Eq. A9 follows from Eq. A2, Eq. A10 follows from the fact that $p(\mathcal{D}_{T},\theta_{S}\,|\,\pi_{\psi},\theta)=p(\mathcal{D}_{T}\,|\,\pi_{\psi})p(\theta_{S}\,|\,\mathcal{D}_{T})$, and Eq. A11 follows from the definition of KL divergence.
@@
Now, we consider the difference between $\text{sEPIG}_{p_{\star}}(\psi)$ and $\mathcal{J}_{p_{\star}}^{{y}^{\star}}(\psi)$:
$\begin{split}\text{sEPIG}_{p_{\star}}(\psi)-\mathcal{J}_{p_{\star}}^{{y}^{\star}}(\psi)&=\mathbb{E}_{p(\theta)p(\mathcal{D}_{T}\,|\,\pi_{\psi},\theta)p_{\star}(x^{\star})p({y}^{\star}\,|\,x^{\star},\theta)}\left[\log\frac{p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}{q_{\phi}({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}\right]\\ &=\mathbb{E}_{p_{\star}(x^{\star})p(\mathcal{D}_{T}\,|\,\pi_{\psi})}\left[\mathbb{E}_{p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}\left[\log\frac{p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}{q_{\phi}({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}\right]\right].\end{split}$ (A13)
@@
$\text{sEPIG}_{p_{\star}}(\psi)-\mathcal{J}_{p_{\star}}^{{y}^{\star}}(\psi)=\mathbb{E}_{p_{\star}(x^{\star})p(\mathcal{D}_{T}\,|\,\pi_{\psi})}[\text{KL}(p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})||q_{\phi}({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T}))].$ (A14)
@@
$\mathcal{J}_{p_{\star}}^{{y}^{\star}}(\psi)\leq\text{sEPIG}_{p_{\star}}(\psi),$ (A15)

### v2

@@ 3.4 Architecture
Aline employs a single, integrated neural architecture based on Transformer Neural Processes (TNPs) [53, 15], a network architecture that has been successfully applied to various amortized sequential decision-making settings [37, 1, 50, 39, 76]. Aline leverages TNPs to concurrently manage historical observations, propose future queries, and condition predictions on specific, potentially varying, inference objectives. An overview of Aline’s architecture is provided in Figure˜2.
The model inputs are structured into three sets. Following standard TNP-based architecture, the context set $\mathcal{D}_{t}=\{(x_{i},y_{i})\}_{i=1}^{t}$ comprises the history of observations, and the target set $\mathcal{T}$ contains the specific target specifier $\xi$. To facilitate active data acquisition, we incorporate a query set $\mathcal{Q}=\{x^{\text{q}}_{n}\}_{n=1}^{N}$ of candidate points. In this paper, we focus on a discrete pool-based setting for consistency, though it can be straightforwardly extended to continuous design spaces (e.g., [64]). Details regarding the embeddings of these inputs are provided in Section˜B.2.
Standard transformer attention mechanisms process these embedded representations. Self-attention operates within the context set, capturing dependencies within $\mathcal{D}_{t}$. Both the query and target set then employ cross-attention to attend to the processed context set representations. To enable the policy $\pi_{\psi}$ to dynamically adapt its acquisition strategy based on the specific inference goal $\xi$, we introduce an additional query-target cross-attention mechanism to allow the query candidates to directly attend to the target set. This allows the evaluation of each candidate $x^{\text{q}}_{n}$ to be informed by its relevance to the different potential targets $\xi$. Examples of the attention mask are shown in Figure˜A1.
Finally, two specialized output heads operate on these processed representations. The inference head ($q_{\phi}$), following ACE [15], uses a Gaussian mixture to parameterize the approximate posteriors and posterior predictives. The acquisition head ($\pi_{\psi}$) generates a policy over the query set $\pi_{\psi}(x_{t+1}|\mathcal{D}_{t})$, drawing on principles from policy-based design methods [37, 50]. This unified design, which leverages a shared transformer backbone with specialized heads, significantly improves parameter efficiency by avoiding the need for separate encoders for the two tasks, leading to faster training and more efficient deployment.
@@ 4 Experiments
We now empirically evaluate Aline’s performance in different active data acquisition and amortized inference tasks. We begin with the active learning task in Section˜4.1, where we want to efficiently minimize the uncertainty over an unknown function by querying $T$ data points. Then, we test Aline’s policy on standard BED benchmarks in Section˜4.2. In Section˜4.3, we demonstrate the benefit of Aline’s flexible targeting feature in a psychometric modeling task [72]. Finally, to demonstrate the scalability of Aline, we test it on a high-dimensional task of actively exploring hyperparameter performance landscapes; the results are presented in Section˜D.1. The code to reproduce our experiments is available at: https://github.com/huangdaolang/aline.
@@ 4.1 Active learning for regression and hyperparameter inference
For the active learning task, Aline is trained on a diverse collection of fully synthetic functions drawn from Gaussian Process (GP) [62] priors (see Section˜C.1.1 for details). We evaluate Aline’s performance under both in-distribution and out-of-distribution settings. For the in-distribution setting, Aline is evaluated on synthetic functions sampled from the same GP prior that is used during training. In the out-of-distribution setting, we evaluate Aline on benchmark functions (Higdon, Goldstein-Price, Ackley, and Gramacy) unseen during training, to assess generalization beyond the training regime. We compare Aline against non-amortized GP models equipped with standard acquisition functions such as Uncertainty Sampling (GP-US), Variance Reduction (GP-VR) [74], EPIG (GP-EPIG) [67], and Random Sampling (GP-RS). Additionally, we include an amortized neural process baseline, the Amortized Conditioning Engine (ACE) [15], paired with Uncertainty Sampling (ACE-US), to specifically evaluate the advantage of Aline’s learned acquisition policy over using a standard acquisition function with an amortized inference model. The performance metric is the root mean squared error (RMSE) of the predictions on a held-out test set.
@@
Results for the active learning task in Figure˜3 show that Aline performs comparably to the best-performing GP-based methods for the in-distribution setting. Importantly, for the out-of-distribution setting, Aline outperforms the baselines in 3 out of the 4 benchmark functions. These results highlight the advantage of Aline’s end-to-end learning strategy, which obviates the need for kernel specification using GPs or explicit acquisition function selection. Further evaluations on additional benchmark functions (Gramacy 1D, Branin, Three Hump Camel), visualizations of Aline’s sequential querying strategy with corresponding predictive updates, and a comparison of average inference times are provided in Section˜D.2.
Additionally, for the in-distribution setting, we test Aline’s capability to infer the underlying GP’s hyperparameters—without retraining, by leveraging Aline’s flexible target specification at runtime. For baselines, we use ACE-US and ACE-RS, since ACE [15] is also capable of posterior estimation. Figure˜4 shows that Aline yields higher log probabilities of the true parameter value under the estimated posterior at each step compared to the baselines. This is due to the ability to flexibly switch Aline’s acquisition strategy to parameter inference, unlike other active learning methods. We also visualize the obtained posteriors in Section˜D.2.
@@ 4.2 Benchmarking on Bayesian experimental design tasks
We test Aline on two classical BED tasks: Location Finding [66] and Constant Elasticity of Substitution (CES) [3], with two- and six-dimensional design space, respectively. As baselines, we include a random design policy, a gradient-based method with variational Prior Contrastive Estimation (VPCE) [25], and three amortized BED methods: Deep Adaptive Design (DAD) [23], vsOED [65], and RL-BOED [9]. Details of the tasks and the baselines are provided in Section˜C.2.
To evaluate performance, we compute a lower bound of the total EIG, namely the sequential Prior Contrastive Estimation lower bound [23]. As shown in Table˜2, Aline surpasses all the baselines in the Location Finding task and achieves competitive performance on the CES task, outperforming most other methods, except RL-BOED. Notably, Aline’s training time is reduced as its reward is based on the internal posterior improvement and does not require a large number of contrastive samples to estimate sEIG. While Aline’s deployment time is slightly higher than MLP-based amortized methods due to the computational cost of its transformer architecture, it remains orders of magnitude faster than non-amortized approaches like VPCE. Visualizations of Aline’s inferred posterior distributions are provided in Section˜D.3.
@@
EIG lower
bound ($\uparrow$)
Training
time (h)
Deployment
time (s)
EIG lower
bound ($\uparrow$)
Training
time (h)
Deployment
time (s)
[…]
@@ 4.3 Psychometric model
Our final experiment involves the psychometric modeling task [72]—a fundamental scenario in behavioral sciences, from neuroscience to clinical settings [29, 57, 73], where the goal is to infer parameters governing an observer’s responses to varying stimulus intensities. The psychometric function used here is characterized by four parameters: threshold, slope, guess rate, and lapse rate; see Section˜C.3.1 for details. Different research questions in psychophysics necessitate focusing on different parameter subsets. For instance, studies on perceptual sensitivity primarily target precise estimation of threshold and slope, while investigations into response biases or attentional phenomena might focus on the guess and lapse rates. This is where Aline’s unique flexible querying strategy can be used to target specific parameter subsets of interest.
We compare Aline with two established adaptive psychophysical methods: QUEST+ [70], which targets all parameters simultaneously, and Psi-marginal [58], which can marginalize over nuisance parameters to focus on a specified subset, a non-amortized gold-standard method for flexible acquisition. We evaluate scenarios targeting either the threshold and slope parameters or the guess and lapse rates. Details of the baselines and the experimental setup are in Section˜C.3.2.
@@
Figure˜5 shows the results. When targeting threshold and slope (Figure˜5a), which are generally easier to infer, Aline achieves results comparable to baselines. When targeting guess and lapse rates (Figure˜5c), QUEST+ performs sub-optimally as its experimental design strategy is dominated by the more readily estimable threshold and slope parameters. In contrast, both Psi-marginal and Aline lead to significantly better inference than QUEST+ when explicitly targeting guess and lapse rates. Moreover, Aline offers a $10\times$ speedup over these non-amortized methods (See Appendix˜D). We also visualize the query strategies adopted by Aline in the two cases: when targeting threshold and slope (Figure˜5b), stimuli are concentrated near the estimated threshold. Conversely, when targeting guess and lapse rates (Figure˜5d), Aline appropriately selects ‘easy’ stimuli at extreme values where mistakes can be more readily attributed to random behavior (governed by lapses and guesses) rather than the discriminative ability of the subject (governed by threshold and slope). To further demonstrate Aline’s runtime flexibility, we conduct two additional investigations detailed in Section˜D.4. First, we show that a single pre-trained ALINE model can dynamically switch its acquisition target mid-experiment. Second, we validate its ability to generalize to novel combinations of targets that were not seen during training, showing the effectiveness of the query-target cross-attention mechanism.
@@ 5 Conclusion
Currently, Aline operates with pre-defined, fixed priors, necessitating re-training for different prior specifications. Future work could explore prior amortization [15, 71], to allow for dynamic prior conditioning. As Aline estimates marginal posteriors, extending this to joint posterior estimation, potentially via autoregressive modeling [11, 35], is a promising direction. Note that, at deployment, we may encounter observations that differ substantially from the training data, leading to degradation in performance. This issue can potentially be tackled by combining Aline with robust approaches such as [22, 36]. Lastly, Aline’s current architecture is tailored to fixed input dimensionalities and discrete design spaces, a common practice with TNPs [53, 50, 76, 37]. Generalizing Aline to be dimension-agnostic [45] and to support continuous experimental designs [64] are valuable avenues for future research.
@@ Acknowledgements
DH, LA and SK were supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI, 359207). The authors acknowledge the research environment provided by ELLIS Institute Finland. LA was also supported by Research Council of Finland grants 358980 and 356498. SK was also supported by the UKRI Turing AI World-Leading Researcher Fellowship, [EP/W002973/1]. AB was supported by the Research Council of Finland grant no. 362534. The authors wish to thank Aalto Science-IT project, and CSC–IT Center for Science, Finland, for the computational and data storage resources provided.
@@
  • In Appendix˜A, we provide detailed derivations and proofs for the theoretical claims made regarding information gain and variational bounds.
  • In Appendix˜B, we present the complete training algorithm and the specifics of the Aline model.
  • In Appendix˜C, we provide comprehensive details for each experimental setup, including task descriptions and baseline implementations.
  • In Appendix˜D, we present additional experimental results, including further visualizations, performance on more benchmarks, and analyses of inference times.
  • In Appendix˜E, we provide an overview of the computational resources and software dependencies for this work.
@@ A.2 Proof of Proposition 1
$\text{sEPIG}(\psi)$ $:=\mathbb{E}_{p_{\star}(x^{\star})p(\mathcal{D}_{T}\,|\,\pi_{\psi})}[\text{H}[p({y}^{\star}\,|\,x^{\star})]-\text{H}[p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})]]$
@@
Here, Eq. A3 follows from conditioning EPIG on the entire trajectory $\mathcal{D}_{T}$ instead of a single data point $y$, Eq. A4 follows from the definition of entropy $H[\cdot]$, and Eq. A5 follows from noting that $p(\mathcal{D}_{T},{y}^{\star}\,|\,\pi_{\psi},x^{\star})=p(\mathcal{D}_{T}\,|\,\pi_{\psi})p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})$. Next, we combine the expectations and express the joint distribution $p(\mathcal{D}_{T},{y}^{\star}\,|\,\pi_{\psi},x^{\star})=\int p(\theta)p(\mathcal{D}_{T}\,|\,\pi_{\psi},\theta)p({y}^{\star}\,|\,x^{\star},\theta)d\theta$, where, following [67], we assume conditional independence between $\mathcal{D}_{T}$ and ${y}^{\star}$ given $\theta$. This yields:
@@ A.3 Proof of Proposition 2
Using the expressions for $\text{sEIG}_{\theta_{S}}$ and $\mathcal{J}_{S}^{\theta}$ from Eq. 7 and Eq. 8, respectively, and noting that $\log q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})=\sum_{l\in S}\log q_{\phi}(\theta_{l}\,|\,\mathcal{D}_{T})$, we can write the expression for $\text{sEIG}_{\theta_{S}}(\psi)-\mathcal{J}_{S}^{\theta}(\psi)$ as:
@@
$=\mathbb{E}_{p(\mathcal{D}_{T},\theta\,|\,\pi_{\psi})}\left[\log\frac{p(\theta_{S}\,|\,\mathcal{D}_{T})}{q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})}\right]$ (A8)
$=\mathbb{E}_{p(\mathcal{D}_{T},\theta_{S}\,|\,\pi_{\psi})}\left[\log\frac{p(\theta_{S}\,|\,\mathcal{D}_{T})}{q_{\phi}(\theta_{S}\,|\,\mathcal{D}_{T})}\right]$ (A9)
@@
Here, Eq. A8 follows from the fact $p(\theta)p(\mathcal{D}_{T}\,|\,\pi_{\psi},\theta)=p(\mathcal{D}_{T},\theta\,|\,\pi_{\psi})$, Eq. A9 follows from Eq. A2, Eq. A10 follows from the fact that $p(\mathcal{D}_{T},\theta_{S}\,|\,\pi_{\psi})=p(\mathcal{D}_{T}\,|\,\pi_{\psi})p(\theta_{S}\,|\,\mathcal{D}_{T})$, and Eq. A11 follows from the definition of KL divergence.
@@
Now, we consider the difference between $\text{sEPIG}(\psi)$ and $\mathcal{J}_{p_{\star}}^{{y}^{\star}}(\psi)$:
$\begin{split}\text{sEPIG}(\psi)-\mathcal{J}_{p_{\star}}^{{y}^{\star}}(\psi)&=\mathbb{E}_{p(\theta)p(\mathcal{D}_{T}\,|\,\pi_{\psi},\theta)p_{\star}(x^{\star})p({y}^{\star}\,|\,x^{\star},\theta)}\left[\log\frac{p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}{q_{\phi}({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}\right]\\ &=\mathbb{E}_{p_{\star}(x^{\star})p(\mathcal{D}_{T}\,|\,\pi_{\psi})}\left[\mathbb{E}_{p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}\left[\log\frac{p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}{q_{\phi}({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})}\right]\right].\end{split}$ (A13)
@@
$\text{sEPIG}(\psi)-\mathcal{J}_{p_{\star}}^{{y}^{\star}}(\psi)=\mathbb{E}_{p_{\star}(x^{\star})p(\mathcal{D}_{T}\,|\,\pi_{\psi})}[\text{KL}(p({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T})||q_{\phi}({y}^{\star}\,|\,x^{\star},\mathcal{D}_{T}))].$ (A14)
@@
$\mathcal{J}_{p_{\star}}^{{y}^{\star}}(\psi)\leq\text{sEPIG}(\psi),$ (A15)